### v1

# Pyramid-Monozone Synergistic Grasping Policy in Dense Clutter

Abstract
Grasping a diverse range of novel objects from dense clutter poses a great challenge to robots because of the occlusion among these objects. In this work, we propose the Pyramid-Monozone Synergistic Grasping Policy (PMSGP) that enables robots to cleverly avoid most occlusions during grasping. Specifically, we initially construct the Pyramid Sequencing Policy (PSP) to sequence each object in the scene into a pyramid structure. By isolating objects layer-by-layer, the grasp candidates will focus on a single layer during each grasp. Then, we devise the Monozone Sampling Policy (MSP) to sample the grasp candidates in the top layer. Through this manner, each grasp will target the topmost object, thereby effectively avoiding most occlusions. We perform more than 7000 real-world grasping among 300 novel objects in dense clutter scenes, demonstrating that PMSGP significantly outperforms seven competitive grasping methods. All grasping videos are available at: https://www.youtube.com/@chenghaoli4532/playlists.
I INTRODUCTION
Grasping a wide range of novel objects reliably can benefit applications in warehousing, manufacturing, retail, and services. However, when these novel objects are densely cluttered, occlusions make this task extremely challenging. As shown in Fig. 1, the object shape incompleteness caused by occlusions can greatly undermine grasping reliability.
One approach to dense clutter grasping is to build a database of grasps on three-dimensional (3D) object models, using performance metrics based on geometry and physics [1, 2], along with stochastic sampling to model uncertainty [3]. This analytical method relies on a perception system to register sensor data with known objects, but it often fails to generalize to various novel objects in dense clutter. To solve this problem, another approach [4, 5] uses machine learning to train function approximators, such as deep neural networks (DNN), to predict the success probability (quality score-based sampling) of grasp candidates from images by leveraging large datasets of empirical successes and failures. However, quality score-based sampling almost does not consider the occlusion problem, which potentially leads to the optimal grasp being predicted on occluded objects prone to collision.
A common solution is to first segment all objects in the scene to create a mask, and then use this mask to guide the sampling of the quality score [6, 7]. This involves evaluating the relation of each instance (object) and determining whether their grasp candidates might result in collisions. But, this method is constrained by the quality score itself, that is, the quality score does not guarantee that every instance has a viable grasp candidate, even if the instance is not occluded. Furthermore, this method typically requires classifying and segmenting all objects in the scene, and these segmented objects may also have incomplete shapes due to occlusion. Thus, this method still does not effectively solve the occlusion problem and essentially trades high costs for limited improvements.
Figure 1: Dense clutter grasping example: Object occlusions are anywhere in this scene (highlighted in green). The objects are often partially visible and closely piled, so grasping them directly can easily fail.
Regarding these problems, we consider that since the robot typically grasps one object at a time, why not identify an unoccluded region before each grasping and only sample within this region, which would not only be accurate but also avoid redundant analysis?
In this paper, we present a novel grasping policy, called the Pyramid-Monozone Synergistic Grasping Policy (PMSGP), which first leverages the Pyramid Sequencing Policy (PSP) to sequence each object in the scene into a pyramid structure, meaning that before each grasp, only the top layer instance is segmented, thus the grasp candidates will focus on this layer. Then, through the Monozone Sampling Policy (MSP), the sampling will be also concentrated on this top-layer instance, more precisely, we do not use the mask to guide the quality score-based sampling, but directly sample these grasp candidates within the mask. This pyramid and monozone operation helps the robot cleverly avoid most occlusions during grasping, and as the grasping progresses, the extent of occlusion gradually decreases, thereby reducing the difficulty of subsequent grasps. Our contributions are as follows:
  1. 1.
    We simplify the complex occlusion problem in dense clutter grasping through first hierarchical, then locally processing ideas.
  2. 2.
    We propose a novel and cost-effective grasping policy based on pyramid sequencing and monozone sampling.
  3. 3.
    We demonstrate that our method far outperforms seven competitive methods among 300 novel objects in various dense clutter scenes. More importantly, even in extremely cluttered scenes with up to 100 novel objects, our method still exhibits high reliability.
II Related Work
While many grasping frameworks exist, this work only focuses on vision-guided 4-DOF grasping with a parallel-jaw gripper. The 4-DOF grasp framework typically performs grasping in a top-down manner, where the robot moves along the $X$, $Y$ and $Z$-axes and rotates only around the $Z$-axis. During grasping, the parallel-jaw gripper will adjust to an appropriate width based on the size of the object perceived by the depth camera. It is mainly divided into analysis-based methods and learning-based methods, as follows.
II-A Analytic-based Grasping Methods
Analytical methods employ mathematical and physical models of geometry, kinematics, and dynamics to compute stable grasps [1, 2, 3]. However, these methods often fail to transfer to a diverse range of novel objects in dense clutter.
II-B Learning-based Grasping Methods
Learning-based methods can generalize to various novel objects in dense clutter, which typically involve training a function approximator, such as DNN, to predict the success probability of grasp candidates from images by leveraging large datasets of empirical successes and failures. As a result, datasets play a crucial role in these methods. One human-labeled dataset is the Cornell Grasping Dataset [8], which contains around 1k RGB-D images and has been widely used to train grasping models based on convolutional neural networks (CNN) [9], such as [10, 11, 12, 13, 14]. However, this dataset is quite small and consists only of single-object images, which limits the dense clutter grasping capabilities of models trained on it.
The Dex-Net series [4, 15, 16, 17, 18] made significant advancements by generating large synthetic datasets that incorporate various dense clutter scenes. Despite these advancements, this approach did not fully resolve the sim-to-real problem. GraspNet [5, 19, 20], in contrast, constructed a real-world dataset featuring 1 billion grasp labels and nearly 100k images with 190 different dense clutter scenes, supporting both 4-DOF and 6D grasping. This dataset enabled remarkable real grasping performance in dense clutter. However, the sampling used in the above methods relies on predicting the success probability (quality score) of the grasp candidates, which almost does not consider the occlusion problem, potentially leading to the optimal grasp being predicted on occluded objects prone to collision.
Recently, [6, 7] proposed to segment all objects in a scene to create a mask and then to use this mask to guide the sampling of quality scores. This process involves evaluating the relation of each object and assessing whether its grasp candidates might result in collisions. However, the effectiveness of this method is also limited by the quality score itself. Specifically, the quality score does not ensure that every object has a viable grasp candidate, even when the object is not occluded. Additionally, this method typically requires classifying and segmenting all objects in the scene, and these segmented objects may still have incomplete shapes due to occlusion. Therefore, this method also fails to effectively address the occlusion issue and essentially exchanges high costs for only marginal improvements. Unlike this method, we first organize each object in the scene into a pyramid structure, segmenting only the topmost object during each grasp. By segmenting layers by layers, most grasp candidates can be restricted to a single layer. We then directly sample grasp candidates for the object in the top layer for each grasp. This allows the robot to cleverly avoid most occlusions during grasping.
III Grasp Configuration
We have previously provided a brief introduction to the 4-DOF grasping framework. Now, we will elaborate on how the grasp configuration is represented in the image coordinate system and its conversion to the robot end effector coordinate system (eye-in-hand grasping). Specifically, we adopt the same grasp configuration in [11], which is composed of parameters $(x,y,w,h,\theta)$ forming a rotated box. Here, $(x,y)$ represents the center of the box, $w$ and $h$ denote the width and height of the box, and $\theta$ represents the angle of the box relative to the horizontal direction.
Since $h$ is used only for visual representation and not in the conversion process, we denote the grasp configurations in the image and robot end effector coordinate systems as $g_{i}(x,y,w,\theta)$ and $g_{r}(x_{r},y_{r},w_{r},\theta_{r})$, respectively. Here, $(x_{r},y_{r},z_{r})$ represents the grasp position in the robot end effector coordinate system, $w_{r}$ is the opening width of the parallel jaw gripper, and $\theta_{r}$ is the rotation angle of the gripper relative to the $Z$ axis. The parameter conversion between $g_{i}$ and $g_{r}$ is described by the following equations:
$\begin{bmatrix}x_{c}\\ y_{c}\\ z_{c}\end{bmatrix}=\begin{bmatrix}f_{x}^{-1}&0&-c_{x}f_{x}^{-1}\\ 0&f_{y}^{-1}&-c_{y}f_{y}^{-1}\\ 0&0&1\end{bmatrix}\begin{bmatrix}x\\ y\\ 1\end{bmatrix}{d}$ (1)
$(x_{r},y_{r},z_{r})=T_{rc}(x_{c},y_{c},z_{c})$ (2)
$(w_{r},\theta_{r})=P_{ri}(w,\theta)$ (3)
Figure 2: The pipline of PMSGP: First, apply Top View Alignment (TVA) to align the depth camera center above the topmost object, and use Cross-prompted Segmentation (CPS) to isolate the object from the scene, obtaining ${M_{r}}$. Next, input both ${M_{r}}$ and the depth image into the grasping model to generate grasp candidates $G$, refining them with Grasp Angle Calibration (GAC) to produce $G^{\prime}$. Finally, sample the optimal grasp $g^{*}$ using Monozone Grasp Analysis (MGA), and further optimize it into $g^{*}_{f}$ as the final grasp.
The conversion process is divided into two parts. The first part involves converting $(x,y)$: using depth information ($d$) and the camera’s intrinsic parameters ($f_{x}$, $f_{y}$ for focal lengths and $c_{x}$, $c_{y}$ for the image center coordinates), we convert $(x,y)$ from the image coordinate system to the camera coordinate system $(x_{c},y_{c},z_{c})$. This is followed by converting $(x_{c},y_{c},z_{c})$ to the robot end effector coordinate system $(x_{r},y_{r},z_{r})$ using the positional transformation relation $T_{rc}$. The second part involves converting $(w,\theta)$ into $(w_{r},\theta_{r})$ using the projection function $P_{ri}$.
The intrinsic parameters and depth information are directly obtained from the depth camera, and $T_{rc}$ is derived from offline eye-in-hand calibration. $T_{rc}$ depends on the relative orientations (excluding rotation) of the $X$, $Y$, and $Z$ axes between the camera coordinate system and the robot end effector coordinate system. Finally, the projection function $P_{ri}$ allows for manual adjustment of the linear relationship between the gripper’s width $w_{r}$ and rotation $\theta_{r}$ relative to the grasp box’s width $w$ and rotation $\theta$.
Once the final grasp pose in the robot end effector coordinate system $(x_{r},y_{r},z_{r},\theta_{r},\theta_{x}^{*},\theta_{y}^{*})$ is obtained, where $\theta_{x}^{*}$ and $\theta_{y}^{*}$ represent the constant rotations relative to the $X$-axis and the $Y$-axis, the gripper will move to the target pose using inverse kinematics and open up to the width $w_{r}$ to grasp.
IV Proposed Method
We propose a novel grasping policy, the Pyramid-Monozone Synergistic Grasping Policy (PMSGP), designed to improve grasping reliability in dense clutter, As illustrated in Fig. 2. PMSGP is composed of two main modules: the Pyramid Sequencing Policy (PSP) and the Monozone Sampling Policy (MSP). The PSP is used for layering objects in cluttered scenes, with submodules for Top View Alignment (TVA) and Cross-prompted Segmentation (CPS). The MSP samples the topmost objects based on the layering results from the PSP, which includes Grasp Angle Calibration (GAC), Monozone Grasp Analysis (MGA), and Optimal Grasp Refinement (OGR). Overall, The MSP is inherently dependent on the PSP, and the submodules within both are linked. Next, we will provide a detailed explanation for them.
IV-A Pyramid Sequencing Policy (PSP)
How to determine an appropriate grasp sequence in dense clutter to decrease the difficulty of grasping? In this context, we introduce the Pyramid Sequencing Policy (PSP), a method that organizes objects within a dense cluttered scene into a pyramid structure, where each layer contains a single object, and the grasp candidate will focus only on the object at the highest layer each time. It is analogous to disassembling a stack of blocks, but in reverse order; rather than building from the ground up, we sequentially remove objects from the top down. This ensures stability and prevents collapse during the disassembly process.
IV-A1 Top View Alignment
In this section, we first align the pixel corresponding to the minimum depth value within the depth camera view (e.g., 1280 $\times$ 720) with the center pixel of this view by moving the robot, we denote the view before and after movement as $V$ and $V^{\prime}$. After the global view alignment, to assist the grasping model analyze local information and minimize the influence of depth camera errors and noises, we refine the alignment process by narrowing the view to 224 $\times$ 224 and performing two additional consecutive alignment operations, the view after this movement is denoted as $V^{\prime\prime}$. The choice of 224 $\times$ 224 is because grasping models typically accept images of this size as input. Finally, we use the Segment Anything Model (SAM) [21] to segment the aligned view to obtain the mask $M_{f}$, focusing only on the object corresponding to the minimum depth value. That is, the center pixel $V^{\prime\prime}(c)$ of $V^{\prime\prime}$ serves as the prompt for this segmentation.
Through this series of steps, PSP not only effectively segments the topmost object in a densely cluttered scene, which consistently appears near the center of the camera view, but also mitigates the issue of partially visible graspable objects caused by input size constraints of the grasping model. It is important to note that this method is not multi-view grasping [11], which is a complex optimization process. In contrast, our method is more direct and efficient.
IV-A2 Cross-prompted Segmentation
Here, we propose the cross-prompt to optimize segmentation in PSP because the single-point prompt is highly unstable in densely cluttered environments. And, this issue is particularly pronounced when the object’s appearance is complex, such as the food packaging, where only part of the object is segmented. So, this limitation adversely impacts subsequent prediction and sampling of grasp candidates.
Specifically, it begins by applying the Sobel operator [22] to extract the edge of the mask $M_{f}$ obtained from the initial single-point prompt segmentation. We then search the two most distant pixels, which we refer to as ${p_{m}}$ and ${p_{m}^{\prime}}$ within this edge. As shown in Eq. 4, $P_{e}$ means the set of all pixels on this edge, and ${P_{e}\times P_{e}}$ represents the Cartesian product of the set $P_{e}$. Next, we calculate the perpendicular line $\perp_{(p_{m},p_{m}^{\prime})}$ connecting ${p_{m}}$ and ${p_{m}^{\prime}}$, and intersecting this perpendicular line with the edges yields another pair of the most distant pixels, which we refer to as ${p_{m_{p}}}$ and ${p_{m_{p}}^{\prime}}$, as shown in Eq. 5 and Fig. 3. (a). These four points are then used as prompts to perform the second segmentation, resulting in ${M_{s}}$. Finally, ${M_{s}}$ is refined by image dilation processing: a depth threshold is applied and pixels from the first prompt serve as initial points for segmentation to produce ${M_{d}}$. By combining ${M_{d}}$ and ${M_{p}}$, we obtain the refined ${M_{r}}$, as shown in Algorithm 1. This method effectively exploits the geometric constraints of the initial mask, thus achieving significant results in the second segmentation, as demonstrated in our ablation studies.
$(p_{m},p_{m}^{\prime})=\underset{\begin{subarray}{c}(p_{i},p_{j})\in P_{e}\times P_{e}\\ i\neq j\end{subarray}}{\arg\max}\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$ (4)
$(p_{m_{p}},p_{m_{p}}^{\prime})=\perp_{(p_{m},p_{m}^{\prime})}\cap{P_{e}}$ (5)
IV-B Monozone Sampling Policy (MSP)
After isolating the topmost object by PSP, we will sample the grasp candidates within this object. Here, we discard quality score-based sampling and propose the Monozone Sampling Policy (MSP) to sample grasp candidates for this object. In other words, we do not use the instance mask to guide the quality score-based sampling, but instead focus directly on sampling these grasp candidates within the mask. Moreover, MSP comprehensively considers all parameters of the grasp candidates (rotation $\theta$, position $(x,y)$, and width $w$) and analyzes the relationship between the grasp candidates, the object being grasped, and the neighboring objects of the object being grasped. Next, we will elaborate on these components.
IV-B1 Grasp Angle Calibration
An inappropriate grasp angle can easily cause the object to slip or fall during grasping due to the uneven force distribution on both sides of the parallel-jaw gripper. Therefore, ensuring that all grasp candidates (denoted as $G$) are at the optimal angle before sampling is essential, as it directly determines the subsequent sampling effectiveness. Here, we use the grasping model in [13] to obtain the grasp candidates. Specifically, we first extract the edge of the instance mask. For each grasp candidate, we rotate them clockwise in 2 degree intervals until they reach 360 degrees. For each rotation $r$, we find four intersection points between the two long sides of the grasp candidate and this edge, that is, $p_{t_{l}}$, $p_{b_{l}}$, $p_{t_{r}}$, and $p_{b_{r}}$.
Figure 3: The visualization of cross-prompted segmentation (a) and grasp angle calibration (b).
Algorithm 1 PSP
1:  Input: Initial view $V$
2:  Output: Refined mask $M_{r}$// Top View Alignment: First conduct global view alignment to get to ${V^{\prime}}$, then conduct local view alignment to get ${V^{\prime\prime}}$, finally segment the object in ${V^{\prime\prime}}$ to get ${M_{f}}$.
3:  ${V^{\prime}}\leftarrow V$, $V^{\prime\prime}\leftarrow V^{\prime}$
4:  ${M_{f}}\leftarrow SAM(V^{\prime\prime}(c))$// Cross-prompted Segmentation: First find four points, then use them to perform segmentation to generate ${M_{s}}$, finally refine ${M_{s}}$ to obtain ${M_{r}}$ based on ${M_{d}}$.
5:  for $(p_{i},p_{j})\in P_{e}\times P_{e},i\neq j$ do
6:     $(p_{m},p_{m}^{\prime})\leftarrow\underset{\begin{subarray}{c}(p_{i},p_{j})\in P_{e}\times P_{e}\\ i\neq j\end{subarray}}{\arg\max}\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$
7:  end for
8:  $(p_{m_{p}},p_{m_{p}}^{\prime})\leftarrow\perp_{(p_{m},p_{m}^{\prime})}\cap{P_{e}}$
9:  ${M_{s}}\leftarrow SAM(p_{m},p^{\prime}_{m},p_{m_{p}},p_{m_{p}}^{\prime})$
10:  Perform dilation to get ${M_{d}}$
11:  ${M_{r}}\leftarrow{M_{s}}+{M_{d}}$
12:  return  ${M_{r}}$
Subsequently, we calculate the angle $\theta^{\prime}$ between the vector $\mathbf{v}_{p_{l}}$ determined by $p_{t_{l}}$ and $p_{b_{l}}$ and the vector $\mathbf{v}_{g_{u}}$ of the long upper side of this grasp candidate, and similarly to get the angle $\theta^{\prime\prime}$ between the vector $\mathbf{v}_{p_{r}}$ determined by $p_{t_{r}}$ and $p_{b_{r}}$ and the vector $\mathbf{v}_{g_{u}}$, shown in Fig. 3. (b). By subtracting 90 degrees from each of these angles, taking the absolute value, and summing them, we obtain the angle difference for each rotation. Finally, we select the rotation with the smallest angle difference ${R}^{*}$(in Eq. 6) and use this rotated box as the new grasp candidate.
[…]
IV-B2 Monozone Grasp Analysis
After grasp angle calibration, we individually analyze each grasp candidate $g_{i}$ to determine the optimal grasp. Specifically, denoting $G^{\prime}$ as grasp candidate sets after angle calibration, we first examine the relationship between each grasp candidate $g_{i}$ and the instance mask $M_{r}$ by checking whether the pixels ${P_{i}}^{s}$ along the two short sides of the grasp candidate fall within the instance mask. So, $g_{i}$ will be filtered out if any pixel $p$ in ${P_{i}}^{s}$ is within mask $M_{r}$. Additionally, to ensure that grasp candidates are concentrated within $M_{r}$, we filter out $g_{i}$ whose centers $c_{i}$ are not within $M_{r}$ and will get grasp candidate sets $G^{\prime\prime}$, this is shown in Eq. 7. However, these processes probable filter all $g_{i}$ out, we adaptively rotate the viewpoint clockwise to alter grasp candidate. That is, if no $g_{i}$ are available from the current viewpoint, we rotate the image 30 degrees at a time and repeat the previous steps until the available grasp candidate are found, as shown in Fig. 4.
Next, based on $G^{\prime\prime}$, we analyze their relationship with adjacent objects by setting a depth threshold $T_{d}$, that is, if the depth difference between any $p$ and $c_{i}$ exceeds $T_{d}$, the grasp candidate $g_{i}$ will be filtered out and get grasp candidate sets $G^{\prime\prime\prime}$, as shown in Eq. 8. So the above steps can effectively filter out most grasp candidates $g_{i}$ that could lead to self-collisions, collisions with adjacent objects, and collisions with the ground.
Finally, based on the sampled grasp candidate sets $G^{\prime\prime\prime}$, we use our previous method [23] to select the $g_{i}$ with the smallest center pixel depth value $d(c_{i})$ as the optimal grasp $g^{*}$, which is shown in Eq. 9. This operation ensures a more stable and safer grasp, especially when dealing with the object with complex shapes.
$G^{\prime\prime}=\{g_{i}\in G^{\prime}\mid({P_{i}}^{s}\cap M_{r}=\emptyset)\land(c_{i}\notin M_{r})\}$ (7)
$G^{\prime\prime\prime}=\{g_{i}\in G^{\prime\prime}\mid\forall p\in{P_{i}}^{s},\,|d(p)-d(c_{i})|\leq T_{d}\}$ (8)
$g^{*}=\mathop{\arg\min}\limits_{g_{i}\in G^{\prime\prime\prime}}d(c_{i})$ (9)
IV-B3 Optimal Grasp Refinement
Although the optimal grasp $g^{*}$ was obtained through the previous coarse-to-fine analysis, this analysis will be influenced by random errors from the depth camera, where $g^{*}$ might still result in collisions with adjacent objects during grasping execution. Spurred from [24], which defines a series of intervals within the grasp box and adjusts the grasp width and position based on the relationships between these intervals. However, this method relies on local regions of the depth image with errors, and it is computationally cumbersome. So, unlike this method, we directly find the minimum rectangle $R_{ec}$ intersecting the optimal grasp $g^{*}$ and the instance mask $M_{r}$, and this process can be represented by $f(g^{*},M_{r})$. Then calculate the shortest width $w_{s}$ and new center point $c^{\prime}_{i}$ of this grasp through ${min{R_{ec}}(w)}$ and ${R_{ec}}(c)$. Additionally, to mitigate the impact of hand-eye calibration errors, we further expand $w_{s}$ to $w^{\prime}_{s}$ by adding some of the hand-eye calibration translation errors $e_{c}$ in the $X$, $Y$-axes. Finally, we use $w^{\prime}_{s}$ and $c^{\prime}_{i}$ as the new width and center of the grasp for optimal grasp, and denote the final grasp as $g^{*}_{f}$. As shown in Eq. 10, Eq. 11, and Algorithm 2.
$R_{ec}=f(g^{*},M_{r})$ (10)
$\left(w^{\prime}_{s},c^{\prime}_{i}\right)=\left(min{R_{ec}}(w)+e_{c},{R_{ec}}(c)\right)$ (11)
Figure 4: The visualization of adaptive viewpoint rotation in MGA
Algorithm 2 MSP
1:  Input: Initial grasp candidate sets $G$
2:  Output: Final grasp $g^{*}_{f}$// Grasp Angle Calibration: Rotate all grasp candidates $g_{i}$ to optimal angle $R^{*}$.
3:  for $g_{i}\in G$ do
[…]
5:     ${G^{\prime}}\leftarrow(G,R^{*})$
6:  end for// Monozone Grasp Analysis: Filter out $g_{i}$ that results in self-collision, collision with adjacent objects, or collision with the ground to obtain the optimal grasp $g^{*}$.
7:  for $g_{i}\in G^{\prime}$ do
8:     $G^{\prime\prime}\leftarrow\{g_{i}\in G^{\prime}\mid({P_{i}}^{s}\cap M_{r}=\emptyset)\land(c_{i}\notin M_{r})\}$
9:  end for
10:  for $g_{i}\in G^{\prime\prime}$ do
11:     $G^{\prime\prime\prime}\leftarrow\{g_{i}\in G^{\prime\prime}\mid\forall p\in{P_{i}}^{s},\,|d(p)-d(c_{i})|\leq T_{d}\}$
12:  end for
13:  for $g_{i}\in G^{\prime\prime\prime}$ do
14:     $g^{*}\leftarrow\mathop{\arg\min}\limits_{g_{i}\in G^{\prime\prime\prime}}d(c_{i})$
15:  end for// Optimal Grasp Refinement: Adjust the width and center of $g^{*}$ to get the final grasp $g^{*}_{f}$.
16:  $R_{ec}\leftarrow f(g^{*},M_{r})$
17:  $\left(w^{\prime}_{s},c^{\prime}_{i}\right)\leftarrow\left(min{R_{ec}}(w)+e_{c},{R_{ec}}(c)\right)$
18:  $g^{*}_{f}\leftarrow(g^{*},(w^{\prime}_{s},c^{\prime}_{i}))$
19:  return  $g^{*}_{f}$
V Experiments
In this section, we validate the effectiveness of our method by conducting extensive experiments. Firstly, we compare it with baseline grasping methods in various mid-clutter (up to 20 objects) and high-clutter scenes(up to 50 objects). Then we extend the cluttered objects to 100 and analyze the effectiveness of PSP and MSP.
V-A Experimental Settings
V-A1 Setting for Grasping Model
The baseline methods we will compare are categorized into two groups. The first group includes GGCNN [10], GGCNN2 [11], GRconvnet [13], SEnet [12], and FCGnet [14], which are suitable for mid-clutter scenarios. The second group comprises DexNet 4.0 [4] and GraspNet [5], which are tailored for high-clutter scenarios. For the first group, since the pre-trained models were all trained on the Cornell Grasping Dataset [8], their performance in cluttered environments is limited. Therefore, we merge the OCID Grasping Dataset [6] with the cluttered scene into the Cornell Grasping Dataset and retrain these models using the parameter settings specified in their original papers (except, all using the RGBD modality). Specifically, we randomly shuffle the entire dataset, using 90% for training and 10% for testing. During training, the data are uniformly cropped to fit the acceptable sizes, the number of training epochs is set to 50, and data augmentation (random zoom and random rotation) is applied. For testing, we use the same metric [11] to report the detection accuracy (Acc) of these methods.
For the second group, we directly use their pre-trained models: the parallel-jaw version of DexNet 4.0 and the planar version of GraspNet [5, 25]. Finally, unless specified, the segmentation and grasp candidate prediction components of PMSGP use the pre-trained models of SAM and GRconvnet in all experiments.
V-A2 Setting for Real Grasping
Our overall grasping system consists primarily of an Intel RealSense D435 depth camera and an Xarm5 robot. We employ an eye-in-hand grasping architecture, with the camera mounted on the robot’s end effector and facing downward. The objects used in our grasping experiments, including 300 novel objects: 100 household objects, 100 snacks, 50 ragdolls, and 50 toys and the grasping difficulty increases successively, as shown in Fig. 5.
Before grasping, we use the method from [10] to fill the depth hole and set a depth value threshold (with the upper limit close to the grasping ground and the lower limit at 10 cm to the camera) to ensure that the grasp is executed within this range. During grasping, each method is tested in five trials per experiment, and the number of failed grasps in each trial (T) is recorded. The grasp success rate (GSR) is calculated by dividing the total number of successful grasps by the total number of grasp attempts across five trials. In addition, to improve experimental safety and ensure all objects are grasped in each trial, we provide minimal manual assistance during the experiments. Specifically, if an object fails to be grasped or recognized 2-3 times, we manually pick up the object and count it as a failure. Additionally, if an object moves out of the camera view, it is repositioned with manual intervention. Similarly, if a grasped object moves out of the robot’s range, causing it to stop, the object is repositioned manually too, and the code will be re-executed.
V-B Comparision Studies
V-B1 Comparision with Baseline Methods in Mid-clutter
In this section, we compare PMSGP with the baseline methods in the first group. We used 10 snacks and 10 household objects to form a mid-clutter scene. The results are shown in I. PMSGP achieves a GSR of 93.5% (100/107), with only 7 total grasp failures, which is far superior to other baselines, except for FCGnet. Additionally, we find that some baselines perform well on the benchmark but not in real grasping. For example, GGCNN2 has a GSR of only 47.6% (100/210) with a total of 110 grasp failures, indicating that this method does not generalize well to novel objects in mid-clutter.
Figure 5: The objects for the grasping experiment. On the first row, the left side are toys, and the right side are ragdolls. On the second row, the left side are snacks, and the right side are household objects.
TABLE I: The GSR between PMSGP and first-group baseline methods
Methods T1 T2 T3 T4 T5 Acc (%) GSR (%)
GGCNN 3 6 5 5 2 22.3 82.6
GGCNN2 16 24 29 23 18 37.7 47.6
GRconvnet 10 3 6 4 5 52.0 78.1
SEnet 8 13 0 5 10 45.0 73.5
FCGnet 4 0 1 3 4 52.0 89.3
PMSGP 1 0 2 1 3 - 93.5
Figure 6: The grasping performance of PMSGP on objects of different grasping difficulties at different scene complexity. The first row is a mid-clutter scene consisting of 20 ragdolls, the second row is a high-clutter scene consisting of 50 snacks, and the last row is an extreme-clutter scene consisting of 100 household objects. The scene complexity and object grasping difficulty of each row increase in sequence.
TABLE II: The GSR between PMSGP and second-group baseline methods
Methods Ragdolls Snacks Toys Households
T1 T2 T3 T4 T5 GSR (%) T1 T2 T3 T4 T5 GSR (%) T1 T2 T3 T4 T5 GSR (%) T1 T2 T3 T4 T5 GSR (%)
DexNet 4.0 4 1 1 3 2 95.8 10 15 7 10 12 82.2 29 23 24 28 30 65.1 29 29 28 26 26 64.4
GraspNet 5 6 3 2 6 92.0 13 10 6 18 14 80.4 25 27 28 21 21 67.2 17 38 30 35 36 61.6
PMSGP 0 1 1 2 0 98.4 4 4 3 3 1 94.3 6 10 10 7 5 86.8 8 6 10 11 5 86.2
V-B2 Comparision with Baseline Methods in High-clutter
DexNet 4.0 and GraspNet are considered state-of-the-art (SOTA) for learning-based 4-DOF and 6D grasping, respectively. Therefore, to demonstrate the effectiveness of PMSGP’s grasping capability, we compare it with these two challenging methods in this section. It is important to note that because of the equipment limitations, we just compare the parallel-jaw version of DexNet 4.0 and the planar version of GraspNet. Moreover, we conducted experiments on high-clutter scenes composed of 50 ragdolls, 50 snacks, 50 toys, and 50 household objects, respectively. The experimental results are shown in II, indicate that PMSGP achieves GSR of 98.4% (250/254) for ragdolls, 94.3% (250/265) for snacks, 86.8% (250/288) for toys, and 86.2% (250/290) for household objects, all of which surpass DexNet 4.0 and GraspNet. More importantly, as the difficulty of the objects being grasped increases, the gap between PMSGP and the baseline methods becomes more obvious. For example, when grasping toys and household objects, PMSGP’s GSR exceeds theirs by up to 20%, demonstrating the high reliability and stability of our method.
V-C Ablation Studies
V-C1 Effectiveness of Pyramid Sequencing Policy
To emphasize the importance of TVA in PSP, we decompose PSP into TVA and Cross-prompted CPS for separate validation. For TVA, we use a mid-clutter scenario consisting of 20 household objects because PMSGP without TVA can only grasp within a very limited area. Additionally, PMSGP without TVA uses the best quality pixel as the prompt for segmentation, with other aspects consistent with PMSGP. For CPS, we use an extreme-clutter scenario consisting of 100 snacks, as their complex appearance helps to fully validate the effectiveness of CPS. Here, PMSGP without CPS uses the first segmentation, with other aspects also consistent with PMSGP too. The experimental results are shown in III and IV. The GSR of PMSGP without TVA is only 74.6% (100/134), whereas the original PMSGP achieves 90% (100/111), a gap of 15.4%, indicating that selecting a reasonable grasping order to avoid occlusions is crucial. Similarly, the GSR of PMSGP without CPS is 79.1% (500/632), compared with the original PMSGP’s 90% (500/564), proving the effectiveness of CPS. In addition, We show some grasping performance of PMSGP on objects of different grasping difficulties at different complexity in Fig 6. It can be seen that PMSGP is always able to avoid most occlusions and focus on the topmost objects for grasping.
V-C2 Effectiveness of Monozone Sampling Policy
In this section, we demonstrate the effectiveness of MSP under the highest grasping difficulty, specifically by using extreme-clutter scenes composed of 100 household objects. These objects exhibit the greatest variation in materials, shapes, and appearances compared to other objects that we used. Additionally, the version without MSP uses quality-based sampling, while other aspects remain consistent with PMSGP. The experimental results, as shown in V, indicate that the GSR of PMSGP without MSP is 75% (500/667), compared to 84.9% (500/589) for PMSGP, a difference of approximately 10%, illustrating the obvious advantage of MSP over quality-based sampling in high-difficulty scenarios.
V-D Failure Case Analysis
In the above experiments, we performed more than 7,000 grasp attempts and achieved a total of 5,800 successful grasps. More importantly, we test PMSGP’s capability in extreme-clutter scenarios involving 100 novel objects. To the best of our knowledge, we have not seen any work that can make it. However, PMSGP still encounters some failures. The first is object slippage, which we plan to address by incorporating suction. The second is that due to the error of the depth camera, some grasps will slightly deviate from the topmost object. This can be achieved by implementing a high-precision industrial depth camera. The final involves objects with similar depths that are tightly packed together. This can be solved by designing a deformable gripper that deforms under force to separate objects.
TABLE III: The impact of with or without TVA
Methods T1 T2 T3 T4 T5 GSR (%)
Without TVA 6 5 10 6 7 74.6
PMSGP 4 1 1 4 1 90.0
TABLE IV: The impact of with or without CPS
Methods T1 T2 T3 T4 T5 GSR (%)
Without CPS 23 29 20 28 32 79.1
PMSGP 19 14 9 9 13 88.7
TABLE V: The impact of with or without MSP
Methods T1 T2 T3 T4 T5 GSR (%)
Without MSP 25 35 32 34 41 75.0
PMSGP 22 14 17 12 24 84.9
VI Conclusion
In this paper, we propose the Pyramid-Monozone Synergistic Grasping Policy that enables robots to effectively avoid most occlusions in dense clutter grasping. By conducting extensive experiments, we demonstrate the effectiveness and reliability of our method. Future work can be divided into two major parts. The first part can focus on addressing the issues highlighted in the Failure Case Analysis to enhance the method proposed in this paper. The second part can involve using this method as a baseline and extending it to human-robot interaction for specific object retrieval. For instance, safely grasping a specific object in a cluttered scene without interfering with other objects, and securely handing it over to a person should be a fascinating point.